OpenStudy (anonymous):

help please with 2 questions

OpenStudy (hoblos):

questions ?

OpenStudy (anonymous):

What is \[\sqrt{21}\] approximated to the nearest tenth?

OpenStudy (anonymous):

Between what two consecutive integers does \[\sqrt{285}\] lie?

OpenStudy (anonymous):

What is \[\sqrt{50}\] approximated to the nearest tenth?

OpenStudy (hoblos):

\[\sqrt{21}= 4.582575\] do you know how to round to the nearest tenth ?

OpenStudy (anonymous):

nope

OpenStudy (hoblos):

the tenth is the number just to the right of the decimal  you must check the number that follows the tenth if it is less then 5, keep the tenth as it is if greater or equal 5, add one to the tenth and all the numbers that follow the tenth the replaced by zeros

OpenStudy (anonymous):

im so confused

OpenStudy (hoblos):

i'll solve this one for you we have the number 4.582575 the tenth is 5 the number that follows it is 8 8>5 so we add one to the tenth 5+1=6 so the number will be 4.6

OpenStudy (anonymous):

so how would you do the other ones...?

OpenStudy (hoblos):

\[\sqrt{50}=7.071\] try to solve this one the same way i did to the previous

OpenStudy (anonymous):

so this one is 7.1...?

OpenStudy (hoblos):

yep you're right

OpenStudy (anonymous):

what about the other one..?

OpenStudy (hoblos):

√285 = 16.882 16<16.882<?

OpenStudy (anonymous):

16.9...?

OpenStudy (hoblos):

integers are whole numbers, they don't contain decimal 1,2,3,4,5,6..

OpenStudy (anonymous):

im confused

OpenStudy (hoblos):

an integer can't contain a decimal (.)